HARRY IS FIVE YEARS OLDER THAN DON. THREE YEARS AGO DON WAS TWO-THIRDS AS OLD AS HARRY. HOW OLD IS HARRY?
I AM NOT SURE BUT I THINK ITS THE 2/3 THATS IS THROWING ME OFF.
THANKS :)

Answer by nerdybill(5404)

HARRY IS FIVE YEARS OLDER THAN DON. THREE YEARS AGO DON WAS TWO-THIRDS AS OLD AS HARRY. HOW OLD IS HARRY?
.
Let x = Don's age
then from "HARRY IS FIVE YEARS OLDER THAN DON." we have:
x+5 = Harry's age
.
Finally, from "THREE YEARS AGO DON WAS TWO-THIRDS AS OLD AS HARRY." we get our equation: (three years ago means current year minus 3)
x-3 = (2/3)(x+5-3)
x-3 = (2/3)(x+2)
3x-9 = 2(x+2)
3x-9 = 2x+4
x-9 = 4
x = 13 years (Don's age)
.
Harry's age:
x+5 = 13+5 = 18 years